We continue to drill Hole U1468B.
There were some problems with the hole yesterday evening at about ~130
The drillers experienced excessive torque and they had to work the pipe
back up to 43.5 mbsf to conduct a mud sweep of the hole.
Once the hole was
reamed and cleaned, drilling continued.In the morning we were at about 700 mbsf,
with the bit advancing at a rate of ~35-40 m/hr.
In the evening we will stop drilling and
start preparing for logging operations. If everything goes as planned, logging
will start early tonight and continue until tomorrow night.